# How Easy is It to Fool Your Multimodal LLMs? An Empirical Analysis on Deceptive Prompts

### v1

@@ Abstract
The remarkable advancements in Multimodal Large Language Models (MLLMs) have not rendered them immune to challenges, particularly in the context of handling deceptive information in prompts, thus producing hallucinated responses under such conditions. To quantitatively assess this vulnerability, we present MAD-Bench, [^1] Short for MultimodAl Deception Benchmark. a carefully curated benchmark that contains 850 test samples divided into 6 categories, such as non-existent objects, count of objects, spatial relationship, and visual confusion. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4V, Gemini-Pro, to open-sourced models, such as LLaVA-1.5 and CogVLM. Empirically, we observe significant performance gaps between GPT-4V and other models; and previous robust instruction-tuned models, such as LRV-Instruction and LLaVA-RLHF, are not effective on this new benchmark. While GPT-4V achieves 75.02% accuracy on MAD-Bench, the accuracy of any other model in our experiments ranges from 5% to 35%. We further propose a remedy that adds an additional paragraph to the deceptive prompts to encourage models to think twice before answering the question. Surprisingly, this simple method can even double the accuracy; however, the absolute numbers are still too low to be satisfactory. We hope MAD-Bench can serve as a valuable benchmark to stimulate further research to enhance models’ resilience against deceptive prompts.
How Easy is It to Fool Your Multimodal LLMs?
An Empirical Analysis on Deceptive Prompts
Yusu Qian, Haotian Zhang, Yinfei Yang, Zhe Gan Apple {yusuqian,haotian.zhang2,yinfeiy,zhe.gan}@apple.com
@@ 1 Introduction
Figure 1: How easy is it to fool your multimodal LLMs? Our study found that multimodal LLMs, such as LLaVA-1.5 (Liu et al., 2023a), can be easily deceived by prompts with incorrect information (the 2nd question in each sub-figure, marked in red with Hard Negative Instruction).
@@
Recent advancements in Multimodal Large Language Models (MLLMs) (Liu et al., 2023b, a; Wang et al., 2023c; You et al., 2024; Bai et al., 2023b; Liu et al., 2024; Zhu et al., 2024), exemplified by models like GPT-4V(ision) (OpenAI, 2023) and Gemini (Team, 2023), mark a significant milestone in the evolution of AI, extending the capabilities of large language models to the realm of visual understanding and interaction.
@@
However, the sophistication of MLLMs brings with it unique challenges, notably, hallucination. Current studies (Liu et al., 2024; Lee et al., 2023; Yin et al., 2023) have been actively exploring solutions to mitigate hallucination, especially when the model tries to generate long responses. However, there still remains a notable gap in the literature: no work has yet been conducted to focus on comprehensively studying the robustness of MLLMs when confronted with deceptive information in the prompts. [^2] LRV-Instruction (Liu et al., 2023a) is the pioneering work in this direction, while we aim to provide a more comprehensive evaluation with hard negative instructions. Please see Section 2 for a more detailed discussion on related work. Our work aims to fill in this gap. This issue is particularly critical, as it pertains to the reliability and trustworthiness of these models in real-world applications (Liu et al., 2023c), and holds substantial importance for the ongoing development and deployment of such AI systems.
@@
To this end, we present [Uncaptioned image] MAD-Bench, a carefully curated benchmark that contains 850 image-prompt pairs spanning across six deception categories, to systematically examine how MLLMs resolve the conflicts when facing inconsistencies between text prompts and images. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4V (OpenAI, 2023), Gemini-Pro (Team, 2023), to open-sourced models, such as LLaVA-1.5 (Liu et al., 2023a) and CogVLM (Wang et al., 2023c). The evaluation is fully automated via the use of GPT-4. Results shed light on how vulnerable MLLMs are in handling deceptive instructions. For example, Figure 1 illustrates how sensitive LLaVA-1.5 (Liu et al., 2023a) is to the factualness of the input prompt and its consistency with the image. When asked “is there a cat in the image?”, LLaVA-1.5 can successfully identify there is no cat; but when prompted with “what color is the cat in the image?”, the model will imagine there is a cat inside. Empirically, we observe that GPT-4V suffers much less when compared with all the other MLLMs; however, the performance is still not ideal (GPT-4V vs. others: 75% vs. 5%-35% accuracy). Further, previous models that aim to mitigate hallucinations, such as LRV-Instruction (Liu et al., 2024) and LLaVA-RLHF (Sun et al., 2023b), are not effective on this new benchmark.
@@
Finally, we provide a simple remedy to boost performance, which was surprisingly found to be effective to double the models’ accuracy. Specifically, we carefully design a system prompt in the form of a long paragraph to be prepended to the existing prompt, to encourage the model to think carefully before answering the question. This simple approach boosts the accuracy of LLaVA-1.5 from 10.42% to 20.56% (similar boosts for other models); however, the absolute numbers are still too low to be satisfactory. Further research is needed to study how to match GPT-4V’s performance (75.02%).
@@ 2 Related Work
MLLM has become an increasingly hot research topic. Early models primarily focused on large-scale image-text pre-training (Wang et al., 2022b, a; Chen et al., 2022, 2023c; Li et al., 2023c; Driess et al., 2023; Huang et al., 2023; Awadalla et al., 2023; Laurençon et al., 2023). Among them, Flamingo (Alayrac et al., 2022) pioneered the integration of a CLIP image encoder with LLMs through gated cross-attention blocks, showcasing emergent multimodal in-context few-shot learning capabilities, via pre-training over millions of image-text pairs and interleaved image-text datasets (Zhu et al., 2023).
On the other hand, recent research has focused on visual instruction tuning (Zhu et al., 2024; Li et al., 2023a; Ye et al., 2023a; Li et al., 2023b; Chen et al., 2023b). Prominent examples include LLaVA(-1.5) (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023), Qwen-VL (Bai et al., 2023a), CogVLM (Wang et al., 2023c), Emu2 (Sun et al., 2023a), SPHINX (Lin et al., 2023), to name a few. Besides text response generation, recent works have also enabled MLLMs for referring and grounding (Peng et al., 2023a; Chen et al., 2023a; You et al., 2024; Wang et al., 2023d), image segmentation (Lai et al., 2023; Zhang et al., 2023), image editing (Fu et al., 2023b), image generation (Koh et al., 2023; Sun et al., 2023a), etc.
The release of proprietary systems like GPT-4V (OpenAI, 2023) and Gemini (Team, 2023) has elevated the research of MLLMs to new heights. Since GPT-4V’s release, researchers have been exploring its capabilities as well as weaknesses (Zhou et al., 2023; Li et al., 2023f; Liu et al., 2023e; Yang et al., 2023; Cui et al., 2023). As MLLMs become stronger, the development of more challenging benchmarks is essential to push the boundaries of what these models can achieve. In this work, we aim to design a new benchmark to evaluate MLLMs’ resilience against deceptive prompts.
Figure 2: Examples of deceptive prompts used in the proposed MAD-Bench with example model responses.
@@
Existing work on mitigating hallucination in LLMs can be roughly divided into two categories: ($i$) prompt engineering (Si et al., 2023; Cheng et al., 2023; Ji et al., 2023; Jones et al., 2023; Mündler et al., 2023; Vu et al., 2023), and ($ii$) model enhancement (Li et al., 2023d; Chuang et al., 2023; Shi et al., 2023; Elaraby et al., 2023; Tian et al., 2024; Qiu et al., 2023; Leng et al., 2023). These studies laid solid foundations for understanding the causes of hallucinations, such as over-reliance on context, or training data biases.
Similarly, hallucination in MLLMs is also growing to be an important research topic (Liu et al., 2024). There are various categories of hallucinations, such as describing objects that are non-existent in the input image, misunderstanding the spatial relationship between objects in the image, and counting objects incorrectly (Liu et al., 2023d). The two main causes of hallucination in MLLMs found in existing work apart from the potential issues with training data include ($i$) limitations in correctly understanding input images, and ($ii$) language model bias (Wang et al., 2023b). Various methods have been proposed to mitigate hallucination in MLLMs (Lee et al., 2023; Yin et al., 2023; Sun et al., 2023b; Wang et al., 2023a; Liu et al., 2024; Zhai et al., 2023; Zhou et al., 2024; Gunjal et al., 2024; Liu et al., 2023b).
Furthermore, various benchmarks have been proposed to evaluate hallucination in MLLMs. Specifically, POPE (Li et al., 2023e), M-HalDetect (Gunjal et al., 2024), and GAVIE (Liu et al., 2024) evaluated object hallucination. HallusionBench (Guan et al., 2023) evaluated both visual and language hallucination. MMHal-Bench (Sun et al., 2023b) evaluated hallucination in more aspects including relations, attributes, environments, etc. Bingo (Cui et al., 2023) studied hallucination in terms of bias and interference in GPT-4V (OpenAI, 2023).
In this work, we aim to study how easy it is to use deceptive prompts that contain information inconsistent with the image to mislead MLLMs to generate responses with hallucination. Note, that we are not the first to study this. A similar model behavior is called “sycophancy” in the LLM literature (Sharma et al., 2023). Fu et al. (2023a) and Liu et al. (2023a) also constructed prompts with deceiving information to test model robustness. Deceptive prompts are termed “negative instructions” in LRV-Instruction (Liu et al., 2023a) and “text-to-image interference” in the Bingo benchmark (Cui et al., 2023). Different from them, we comprehensively study MLLMs’ ability to handle deceptive prompts in multiple categories. Unlike previous studies (Fu et al., 2023a; Liu et al., 2023a) which primarily used “Is/Are/Can” questions, we found that it is relatively easy for state-of-the-art MLLMs to counter deceptive information in such formats. Consequently, we shifted our focus to questions beginning with “What”, “How”, “Where”, etc., to provide a more challenging and insightful evaluation.
@@ 3 MAD-Bench
In this section, we present MAD-Bench, introduce how we collect [^3] The authors are responsible for collecting the data, Apple Inc is not. deceptive image-prompt pairs, as well as our evaluation method.
@@ 3.1 Deception Categories
MAD-Bench encompasses six distinct categories of 850 image-prompt pairs designed to test the resilience of MLLMs against deceptive prompts. Table 1 provides the statistics of each category, and Figure 2 shows examples of deceptive prompts. The selected categories are partly inspired by Liu et al. (2023d). Below, we detail each category.
Figure 3: Examples of image-prompt pairs in the Visual Confusion category of MAD-Bench.
Count of Object. This category intentionally cites an incorrect quantity of visible objects in the image. A response fails this test if it asserts the presence of $m$ instances of an object ‘A’ when, in reality, a different number $n$ of object ‘A’ is present — $n$ being distinct from $m$ and not zero. The images for this and the subsequent four categories are sourced from COCO 2017 (Lin et al., 2015). Using a public dataset sometimes brings concerns about data leakage. In our case, given the special nature of our deceptive prompts to be introduced in the next section, this will not be a problem. An accurate response would either challenge the prompt’s inconsistency with the visual data and abstain from speculating on absent information, or seek further clarification to resolve any uncertainties.
Deception Category Count Image Source
Count of Object 188 COCO 2017
Non-existent Object 244 COCO 2017
Object Attribute 136 COCO 2017
Scene Understanding 122 COCO 2017
Spatial Relationship 132 COCO 2017
Visual Confusion 28 In the Wild
Table 1: Statistics of the 850 image-prompt pairs in MAD-Bench.
@@
Figure 4: Illustration of the process of generating deceptive prompts in the non-existent object category using GPT-4 and COCO ground-truth captions.
@@
Spatial Relationship. This category presents prompts that incorrectly specify the spatial dynamics between objects that do indeed exist within the image. A misstep in this category arises when a response correctly recognizes the objects but misrepresents their spatial relations.
Visual Confusion. This category is different from the previous ones by employing both the prompts and the images as instruments of deception, often deceptive even to the human eye. This category includes three types of images: ($i$) those depicting naked-eye 3D paintings or screens, ($ii$) visual dislocation photography, and ($iii$) mirror reflections. Figure 3 shows an example image-prompt pair  [^4] Photo credit to Braga last1 and Tiago Silva. in each category. Here, the prompts paired with the 3D paintings or screens aim to deceive the MLLMs by portraying the objects in the two-dimensional artwork as three-dimensional entities. With visual dislocation photography, the prompts reinforce the optical illusions present in the images. Lastly, the prompts associated with mirror reflections attempt to deceive the MLLMs into interpreting reflections as tangible objects.
@@ 3.2 Prompt Generation Method
The process of creating deceptive prompts was automated by employing GPT-4, leveraging the ground-truth captions from the COCO dataset (Lin et al., 2015). We chose not to use GPT-4V for this task, as we later also evaluated GPT-4V on this benchmark, and empirically, employing GPT-4 is already enough for this task. To guide GPT-4 in generating questions that would intentionally mislead MLLMs within the specified categories, we crafted tailored prompts. These guiding prompts are provided in Appendix A.2, from Figure 18 to 20. The process is illustrated in Figure 4, using an example in the non-existent object category. Bounding box information is not used as part of the prompt sent to GPT-4, as empirically, we observed that it does not contribute to further improving the quality of generated prompts in our deceptive categories. Following the generation of these deceptive questions, a rigorous manual filtering process is followed to ensure that each question adheres to its category’s deceptive criteria and maintains relevance to its associated image.
@@ 3.3 Response Evaluation Method
We use GPT-4 to evaluate generated responses from 10 models, including ($i$) 6 open-sourced models: LLaVA-1.5 (Liu et al., 2023a), InstructBLIP (Dai et al., 2023), Ferret (You et al., 2024), Kosmos-2 (Peng et al., 2023b), mPLUG-Owl2 (Ye et al., 2023b), and CogVLM (Wang et al., 2023c), ($ii$) 2 additional open-sourced models that aim to reduce hallucination: LLaVA-RLHF (Sun et al., 2023b) and LRV-V1 (Liu et al., 2024), and ($iii$) 2 state-of-the-art proprietary systems: Gemini-Pro (Team, 2023) and GPT-4V (OpenAI, 2023).
The number of images in the Visual Confusion category is relatively small, while most of them contain humans, so we did not evaluate Gemini in this category as it cannot generate responses for images containing humans. The effect of this on other categories is neglectable. Mirroring the prompt generation method, we design specific prompts for each deceptive category to critically assess the responses. Our primary metric of evaluation is binary, focused strictly on whether the response has been misled, without considering other qualitative aspects such as helpfulness. These prompts for model evaluation are provided in Appendix A.3.
Figure 5: Example failure cases of GPT-4V (OpenAI, 2023).
To verify the accuracy of GPT-4’s automated evaluation, we randomly select 500 responses spanning the various models and deceptive categories for a manual accuracy check. This validation process yielded a 97.0% concordance rate with the outcomes of human evaluation, underlining the reliability of our approach.

### v2

@@ Abstract
The remarkable advancements in Multimodal Large Language Models (MLLMs) have not rendered them immune to challenges, particularly in the context of handling deceptive information in prompts, thus producing hallucinated responses under such conditions. To quantitatively assess this vulnerability, we present MAD-Bench, [^1] Short for MultimodAl Deception Benchmark. a carefully curated benchmark that contains 1000 test samples divided into 5 categories, such as non-existent objects, count of objects, and spatial relationship. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4v, Reka, Gemini-Pro, to open-sourced models, such as LLaVA-NeXT and MiniCPM-Llama3. Empirically, we observe significant performance gaps between GPT-4o and other models; and previous robust instruction-tuned models are not effective on this new benchmark. While GPT-4o achieves 82.82% accuracy on MAD-Bench, the accuracy of any other model in our experiments ranges from 9% to 50%. We further propose a remedy that adds an additional paragraph to the deceptive prompts to encourage models to think twice before answering the question. Surprisingly, this simple method can even double the accuracy; however, the absolute numbers are still too low to be satisfactory. We hope MAD-Bench can serve as a valuable benchmark to stimulate further research to enhance models’ resilience against deceptive prompts.
Figure 1: How easy is it to fool your multimodal LLMs? Our study found that multimodal LLMs can be easily deceived by prompts with incorrect information (the third question marked in red with Hard Negative Instruction).
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) [1, 2, 3, 4, 5, 6, 7], exemplified by models like GPT-4V(ision) [8] and Gemini [9], mark a significant milestone in the evolution of AI, extending the capabilities of large language models to the realm of visual understanding and interaction.
@@
However, the sophistication of MLLMs brings with it unique challenges, notably, hallucination. Current studies [6, 10, 11] have been actively exploring solutions to mitigate hallucination, especially when the model tries to generate long responses. However, there still remains a notable gap in the literature: no work has yet been conducted to focus on comprehensively studying the robustness of MLLMs when confronted with deceptive information in the prompts. [^2] LRV-Instruction [2] is the pioneering work in this direction, while we aim to provide a more comprehensive evaluation with hard negative instructions. Please see Section 2 for a more detailed discussion on related work. Our work aims to fill in this gap. This issue is particularly critical, as it pertains to the reliability and trustworthiness of these models in real-world applications [12], and holds substantial importance for the ongoing development and deployment of such AI systems.
@@
To this end, we present MAD-Bench, a carefully curated benchmark that contains 1000 image-prompt pairs spanning across five deception categories, to systematically examine how MLLMs resolve the conflicts when facing inconsistencies between text prompts and images. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4V [8], Gemini-Pro [9], to open-sourced models, such as LLaVA-NeXT [13] and MiniCPM [14]. The evaluation is fully automated via the use of GPT-4o [15]. Results shed light on how vulnerable MLLMs are in handling deceptive instructions. For example, Figure 1 illustrates how sensitive LLaVA-1.5 [2] is to the factualness of the input prompt and its consistency with the image. When asked “is there a cat in the image?”, LLaVA-1.5 can successfully identify there is no cat; but when prompted with “what color is the cat in the image?”, the model will imagine there is a cat inside. Empirically, we observe that GPT-4V suffers much less when compared with all the other MLLMs; however, the performance is still not ideal (GPT-4V vs. others: 82% vs. mostly 3%-50% accuracy).
@@
Finally, we provide a simple remedy to boost performance, which was surprisingly found to be effective to double the models’ accuracy. Specifically, we carefully design a system prompt in the form of a long paragraph to be prepended to the existing prompt, to encourage the model to think carefully before answering the question. This simple approach boosts the accuracy of LLaVA-NeXT-13b from 49.65% to 68.21% (similar boosts for other models); however, the absolute numbers still have room for improvement.
@@ 2 Related Work
MLLM has become an increasingly hot research topic. Early models primarily focused on large-scale image-text pre-training [16, 17, 18, 19, 20, 21, 22, 23, 24]. Among them, Flamingo [25] pioneered the integration of a CLIP image encoder with LLMs through gated cross-attention blocks, showcasing emergent multimodal in-context few-shot learning capabilities, via pre-training over millions of image-text pairs and interleaved image-text datasets [26].
On the other hand, recent research has focused on visual instruction tuning [7, 27, 28, 29, 30]. Prominent examples include LLaVA(-1.5) [1, 2], InstructBLIP [31], Qwen-VL [32], CogVLM [3], Emu2 [33], SPHINX [34], to name a few. Besides text response generation, recent works have also enabled MLLMs for referring and grounding [35, 36, 4, 37], image segmentation [38, 39], image editing [40], image generation [41, 33], etc.
The release of proprietary systems like GPT-4V [8] and Gemini [9] has elevated the research of MLLMs to new heights. Since GPT-4V’s release, researchers have been exploring its capabilities as well as weaknesses [42, 43, 44, 45, 46]. As MLLMs become stronger, the development of more challenging benchmarks is essential to push the boundaries of what these models can achieve. In this work, we aim to design a new benchmark to evaluate MLLMs’ resilience against deceptive prompts.
@@
Existing work on mitigating hallucination in LLMs can be roughly divided into two categories: ($i$) prompt engineering [47, 48, 49, 50, 51, 52, 53], and ($ii$) model enhancement [54, 55, 56, 57, 58, 59, 60, 61, 62]. These studies laid solid foundations for understanding the causes of hallucinations, such as over-reliance on context, or training data biases.
Similarly, hallucination in MLLMs is also growing to be an important research topic [6]. There are various categories of hallucinations, such as describing objects that are non-existent in the input image, misunderstanding the spatial relationship between objects in the image, and counting objects incorrectly [63]. The two main causes of hallucination in MLLMs found in existing work apart from the potential issues with training data include ($i$) limitations in correctly understanding input images, and ($ii$) language model bias [64]. Various methods have been proposed to mitigate hallucination in MLLMs [1, 6, 10, 11, 65, 66, 67, 68, 69].
Furthermore, various benchmarks have been proposed to evaluate hallucination in MLLMs. Specifically, POPE [70], M-HalDetect [69], GAVIE [6], and Throne [71] evaluated object hallucination. HallusionBench [72] evaluated both visual and language hallucination. MMHal-Bench [65] evaluated hallucination in more aspects including relations, attributes, environments, etc. Bingo [46] studied hallucination in terms of bias and interference in GPT-4V [8]. Hal-Eval [73] assesses event hallucination, which involves creating a fictional target and constructing an entire narrative around it, encompassing its attributes, relationships, and actions.
In this work, we aim to study how easy it is to use deceptive prompts that contain information inconsistent with the image to mislead MLLMs to generate responses with hallucination. Note, that we are not the first to study this. A similar model behavior is called “sycophancy” in the LLM literature [74]. MME [75] and LLaVA-Bench (in-the-Wild) [2] also constructed prompts with deceiving information to test model robustness. Deceptive prompts are termed “negative instructions” in LRV-Instruction [2] and “text-to-image interference” in the Bingo benchmark [46]. Different from them, we comprehensively study MLLMs’ ability to handle deceptive prompts in multiple categories. Unlike previous studies [2, 75] which primarily used “Is/Are/Can” questions, we found that it is relatively easy for state-of-the-art MLLMs to counter deceptive information in such formats. Consequently, we shifted our focus to questions beginning with “What”, “How”, “Where”, etc., to provide a more challenging and insightful evaluation.
@@ 3 MAD-Bench
In this section, we present MAD-Bench, introduce how we collect deceptive image-prompt pairs, as well as our evaluation method. The images in MAD-Bench are sourced from COCO 2017 validation set [76], SBU [77], and TextVQA [78]. Using a public dataset sometimes brings concerns about data leakage. In our case, given the special nature of our deceptive prompts to be introduced in the following section, this will not be a problem.
Figure 2: Examples of deceptive prompts with example model responses.
@@ 3.1 Deception Categories
MAD-Bench encompasses five distinct categories of 1000 image-prompt pairs designed to test the resilience of MLLMs against deceptive prompts.
Deception Category Count
Count of Object 32
Non-existent Object 778
Object Attribute 24
Scene Understanding 115
Text Recognition 51
Table 1: Statistics of the 1000 image-prompt pairs in MAD-Bench.
Table 1 provides the statistics of each category, and Figure 2 shows examples of deceptive prompts. The selected categories are partly inspired by MMBench [63]. Below, we detail each category.
Count of Object. This category intentionally cites an incorrect quantity of visible objects in the image. A response fails this test if it asserts the presence of $m$ instances of an object ‘A’ when, in reality, a different number $n$ of object ‘A’ is present — $n$ being distinct from $m$ and not zero. An accurate response would either challenge the prompt’s inconsistency with the visual data and abstain from speculating on absent information, or seek further clarification to resolve any uncertainties.
@@
Text Recognition. This category presents prompts that incorrectly identifies tech-rich objects in the image as something else or misunderstands the information conveyed in a piece of text. A misstep in this category occurs when a response fails to accurately identify the true information from the text.
@@ 3.2 Prompt Generation Method
The process of creating deceptive prompts was automated by employing GPT-4o. To guide GPT-4o in generating questions that would intentionally mislead MLLMs, we using the following prompt:
Following the generation of these deceptive questions, a rigorous manual filtering process is followed to ensure that each question adheres to its category’s deceptive criteria and maintains relevance to its associated image.
@@ 3.3 Response Evaluation Method
We use GPT-4o to evaluate generated responses from 19 models, including ($i$) 15 open-sourced models: Ferret [4], Kosmos2 [35], CogVLM [3], Yi-VL-34b [79], mPLUG-Owl2  [80], MiniCPM-Llama3-v2.5 [14], Phi-3-vision [81], XComposer2 [82], LLaVA-Next-7b [13], LLaVA-NeXT-13b-vicuna [13], LLaVA-NeXT-34b [13], DeepSeek-VL-7b [83], Idefics-2 [84], Qwen-VL-Chat [5], and InternVL-Chat-v1.5 [85] ($ii$) 4 state-of-the-art proprietary systems: Gemini-Pro [9], Reka [86], GPT-4V [8], and GPT-4o [15].
Mirroring the prompt generation method, we design specific prompts for each deceptive category to critically assess the responses. Our primary metric of evaluation is binary, focused strictly on whether the response has been misled, without considering other qualitative aspects such as helpfulness. These prompts for model evaluation are provided in Appendix.
To verify the accuracy of GPT-4o’s automated evaluation, we randomly select 500 responses spanning the various models and deceptive categories for a manual accuracy check. This validation process yielded a 98.0% concordance rate with the outcomes of human evaluation, underlining the reliability of our approach.